OpenStudy (emilyjones284):

Very unsure on this one... please help. ill give medal and fan (: posting below.

OpenStudy (emilyjones284):

Suppose Q and R are independent events. Find P(Q and R) if P(Q) = \[\frac{ 7 }{ 15}\] and P(R) = \[\frac{ 4 }{ 7 }\]

OpenStudy (queelius):

In general, the probability of events Q and R happening is P(Q|R) * P(R). If Q and R are independent, the P(Q|R) = P(Q). So, with this independence assumption, we simply have: P(Q & R) = P(Q) * P(R)

OpenStudy (emilyjones284):

so i add 0.32 and 0.19?

OpenStudy (queelius):

No, don't add; multiply. In words, the probability of Q and R happening is the probability of Q happening times the probability of R happening.

OpenStudy (emilyjones284):

so multiply 0.32 and 0.19 and then do what?

jimthompson5910 (jim_thompson5910):

how are you getting 0.32 and 0.19 ?

OpenStudy (queelius):

Oh yeah, that's a good point. Those probabilities for Q and R do not correspond to the probabilities given in your question. For instance, P(Q) is given as 7/15.

OpenStudy (emilyjones284):

oops sorry i was looking at a similar question with different values

OpenStudy (queelius):

Ahhh, then yes, you have the right idea, simply multiply them together.

OpenStudy (emilyjones284):

and after i multiply them what do i do?

OpenStudy (queelius):

@jim_thompson5910 Thanks for noticing that!

jimthompson5910 (jim_thompson5910):

np and keep things as fractions to avoid roundoff error

OpenStudy (queelius):

After multiplying them, then you are done. P(Q & R) = P(Q) * P(R)

OpenStudy (emilyjones284):

Ohhh okay thats easy thanks (:

OpenStudy (queelius):

And, as @jim_thompson5910 stated, multiply them as fractions. There is no need to convert them to an approximate decimal form. If you'd like, you can try to do that here.

OpenStudy (emilyjones284):

Okay I got 4/15 (:

OpenStudy (queelius):

Good work! :)